OpenStudy (josh.darnit10):

What can you multiply to get -96 but add to get 14?

OpenStudy (studygurl14):

Set up two equations: xy = -96 x + y = 14

OpenStudy (shadowlegendx):

@josh.darnit10 Do you know how to solve for a system of equations?

OpenStudy (shadowlegendx):

???

OpenStudy (josh.darnit10):

No, @ShadowLegendX

OpenStudy (studygurl14):

Rearrange the second equation and plug into the first x + y = 14 y = 14 - x xy = -96 x(14 - x) = -96 Can you do the rest?

OpenStudy (naka354):

x^2 - 14x - 96 = 0 etc